# Ihara’s Lemma for $\operatorname{GL}_{d}$: the limit case

### v4

@@ 2.3. Kottwiz–Harris–Taylor Shimura varieties
Important property: over ¯$\overline{\mathbb{Z}}_{l}$, there are at least two notions of intermediate extension associated to the two classical $t$-structures $p$ and $p+$. By proposition 2.4.1 of [Boy23b], in the limit case where all $\overline{\mathbb{F}}_{l}$-cuspidal representations are supercuspidal, as recalled after proposition 2.1.3, all the $p$ and $p+$ intermediate extensions of Harris-Taylor local systems coincide.
@@ 3.1. Filtrations of stratification of $\Psi_{\varrho}$
The graded parts $\operatorname{gr}^{t}_{!}(\Psi_{K,\varrho})$ are, by construction, free and admit a strict888meaning the graded parts are free filtration, cf. [Boy14] corollary 3.4.5
@@
Concerning the ¯$\overline{\mathbb{Z}}_{l}$-structures, cf. the third global result of the introduction of [Boy23b] , for every $1\leq k\leq i\leq s$, we have strict epimorphisms999strict means that the cokernel is torsion free
@@
For a finite set $S$ of places of ${\mathbb{Q}}$ containing the places where $G$ is ramified, denote by ${\mathbb{T}}^{S}_{abs}:=\prod^{\prime}_{x\not\in S}{\mathbb{T}}_{x,abs}$ the abstract unramified Hecke algebra where ${\mathbb{T}}_{x,abs}\cong\overline{\mathbb{Z}}_{l}[X^{un}(T_{x})]^{W_{x}}$ for $T_{x}$ a split torus, $W_{x}$ the spherical Weyl group and $X^{un}(T_{x})$ the set of ¯$\overline{\mathbb{Z}}_{l}$-unramified characters of $T_{x}$.
@@
${\mathbb{T}}_{w|_{{\mathbb{Q}}},abs}=\overline{\mathbb{Z}}_{l}\bigl{[}T_{w^{\prime},i}:~{}i=1,\cdots,d,~{}w^{\prime}|(w|_{{\mathbb{Q}}})\bigr{]}$
@@
Consider now the filtration of stratification of $\Psi_{\varrho}:=\Psi_{K^{v}(\infty),\varrho}$131313i.e. with infinite level at $v$ constructed using the adjunction morphisms $j^{=h}_{!}j^{=h,*}$ as in [Boy14]
@@ 3.3. Local and global monodromy
4. Proof of the Ihara’s lemma
@@ 4.2. Level raising
Fact from [Boy10] §3: for an irreducible cuspidal representation $\pi_{v}$ such that its modulo $l$ reduction is isomorphic to $\varrho$, $H^{0}(\operatorname{Sh}_{K^{v}(\infty),\bar{s}_{v}},P(t,\pi_{v}))_{\mathfrak{m}}\otimes_{\overline{\mathbb{Z}}_{l}}\overline{\mathbb{Q}}_{l}$ is the sum of the contributions of $\Pi_{\widetilde{\mathfrak{m}}}$ with $\widetilde{\mathfrak{m}}\subseteq\mathfrak{m}$ such that $\Pi_{\widetilde{\mathfrak{m}}}$ is of the following shape: $\mathrm{st}_{t}(\pi^{\prime}_{v})\times?$ where $\pi^{\prime}_{v}$ is an unramified twist of $\pi_{v}$ and ???? is any representation of $\operatorname{GL}_{d-tg}(F_{v})$ whose cuspidal support is not linked to those of $\mathrm{st}_{t}(\pi^{\prime}_{v})$.

### v5

@@ 2.3. Kottwiz–Harris–Taylor Shimura varieties
Important property: over ¯$\overline{\mathbb{Z}}_{l}$, there are at least two notions of intermediate extension associated to the two classical $t$-structures $p$ and $p+$. By proposition 2.4.1 of [Boy23b], in the limit case where all $\overline{\mathbb{F}}_{l}$-cuspidal representations are supercuspidal, as recalled after proposition 2.1.4, all the $p$ and $p+$ intermediate extensions of Harris-Taylor local systems coincide. The arguments in loc. cit. are rather difficult but if one accepts to restrict to the case where the irreducible constituants of $\overline{\rho}_{\mathfrak{m}}$ are all characters, then the proof of this fact is easy. Indeed as $\mathrm{Sh}^{\geq h}_{K^{v}(n),\bar{s}_{v},1}$ is smooth over $\operatorname{Spec}\overline{\mathbb{F}}_{p}$, then $HT(\chi_{v},\Pi_{h})_{1}$ is perverse for the two $t$-structures with
[…]
@@ 3.1. Filtrations of stratification of $\Psi_{\varrho}$
The graded parts $\operatorname{gr}^{t}_{!}(\Psi_{K,\varrho})$ are, by construction, free and admit a strict999meaning the graded parts are free filtration, cf. [Boy14] corollary 3.4.5
@@
Concerning the ¯$\overline{\mathbb{Z}}_{l}$-structures, cf. the third global result of the introduction of [Boy23b] , for every $1\leq k\leq i\leq s$, we have strict epimorphisms101010strict means that the cokernel is torsion free
@@
For a finite set $S$ of places of ${\mathbb{Q}}$ containing the places where $G$ is ramified, denote by ${\mathbb{T}}^{S}_{abs}:=\prod^{\prime}_{x\not\in S}{\mathbb{T}}_{x,abs}$ the abstract unramified Hecke algebra where ${\mathbb{T}}_{x,abs}\cong{\mathbb{Z}}_{l}[X^{un}(T_{x})]^{W_{x}}$ for $T_{x}$ a split torus, $W_{x}$ the spherical Weyl group and $X^{un}(T_{x})$ the set of ¯$\overline{\mathbb{Z}}_{l}$-unramified characters of $T_{x}$.
@@
${\mathbb{T}}_{w|_{{\mathbb{Q}}},abs}={\mathbb{Z}}_{l}\bigl{[}T_{w^{\prime},i}:~{}i=1,\cdots,d,~{}w^{\prime}|(w|_{{\mathbb{Q}}})\bigr{]}$
@@
Consider now the filtration of stratification of $\Psi_{\varrho}:=\Psi_{K^{v}(\infty),\varrho}$141414i.e. with infinite level at $v$ constructed using the adjunction morphisms $j^{=h}_{!}j^{=h,*}$ as in [Boy14]
@@ 3.3. Local and global monodromy
4. Uniformity of automorphic sub-lattices
4.1. Infinite level at $v$
For $K\in{\mathcal{K}}$ a compact open subgroup of $G({\mathbb{A}}_{\mathbb{Q}}^{\infty})$ as before, we consider the category ${\mathcal{C}}(K)$ of finitely generated ${\mathcal{O}}_{L}$-modules with a continuous $K$-action and let ${\mathcal{C}}^{\prime}(K)$ be a Serre subcategory of ${\mathcal{C}}(K)$. Let ${\mathcal{C}}(K)_{Z}$ be the subcategory of ${\mathcal{C}}(K)$ consisting of those $\sigma\in{\mathcal{C}}$ possessing a central character which lifts $\prod_{v\in\Sigma^{+}(K)}(\det\overline{\rho}_{|I_{v}}\bar{\varepsilon})\circ\operatorname{Art}$ and let ${\mathcal{C}}^{\prime}(K)_{Z}$ be a Serre subcategory of ${\mathcal{C}}(K)_{Z}$.
For each $v\in\Sigma^{+}(K)$, we fix
  • a lift of a geometric Frobenius element $\operatorname{Frob}_{v}\in G_{v}:=\operatorname{Gal}(\overline{F}_{v}/F_{v})$,
  • an element $\alpha_{v}\in\overline{\mathbb{Z}}_{l}$ lifting $\det\overline{\rho}_{v}(\operatorname{Frob}_{v})$,
  • a character $\psi_{\sigma,v}:G_{v}\longrightarrow\overline{\mathbb{Z}}_{l}^{\times}$ such that $\psi_{\sigma,v}(\operatorname{Frob}_{v})=\alpha_{v}$ and the composite with Artin map has a restriction to $I_{v}$ equal to the central character $\sigma$.
For $\sigma\in{\mathcal{C}}^{\prime}(K)_{Z}$, we define
$M_{K}(\sigma):=H^{0}(\mathrm{Sh}_{K,\bar{s}_{v}},\Psi_{v}\otimes\mathcal{L}_{\sigma^{\vee}})_{\mathfrak{m}}^{\vee},$
where $\mathcal{L}_{\sigma^{\vee}}$ is the sheaf associated to $\sigma^{\vee}$. When we consider only the direct factor $\Psi_{\varrho}$ of $\Psi_{v}$, we then denote by $M_{K,\varrho}(\sigma)$ the associated space.
Remark. As the $\mathfrak{m}$-localized cohomology groups are concentrated in middle degree, note that the functor $\sigma\mapsto M_{\infty,K,\varrho}(\sigma)$ is exact.
Let $\sigma^{v}$ be a continuous finitely generated representation of $K^{v}$ and let $\sigma_{0}(v)$ be the representation of $K=K^{v}K_{v}$ given by twisting the action of $K^{v}$ on $\sigma^{v}$ by the character $\psi\circ\operatorname{Art}_{F_{v}}\circ\det$ of $K_{v}$. We assume that $\sigma\in{\mathcal{C}}(K)_{Z}$. We then define
$S_{K^{v}}(\sigma^{v}):=\lim_{\rightarrow n}H^{d-1}(\operatorname{Sh}_{K^{v}(n),\bar{s}_{v}},{\mathcal{L}}_{\sigma_{0}(v)^{\vee})}^{\vee})_{\mathfrak{m}}.$
We let ${\mathbb{T}}(\sigma^{v})$ denote the image of ${\mathbb{T}}^{S}$ in $\operatorname{End}_{\mathcal{O}}(S_{K^{v}}(\sigma^{v}))$ and we write $\rho(\sigma^{v})_{\mathfrak{m}}$ for the composite $\operatorname{Gal}_{F,s}\smash{\mathop{\longrightarrow}\limits^{\rho^{univ}}}\operatorname{GL}_{d}(R_{S}^{univ})\longrightarrow\operatorname{GL}_{d}({\mathbb{T}}(\sigma^{v})_{\mathfrak{m}})$. We set
$M_{K^{v}}(\sigma^{v})\stackrel{{\scriptstyle\textrm{\tiny{{def}}}}}{{=}}\Bigl{(}\hom_{{\mathbb{T}}(\sigma^{v})_{\mathfrak{m}}[\operatorname{Gal}_{F,S}]}(\rho(\sigma^{v})_{\mathfrak{m}},S_{K^{v}}(\sigma_{v})_{\mathfrak{m}}))^{*}.$ (21)
For a fixed $K_{v}$ and a representation $\sigma_{v}$ of $K_{v}$ such that $\sigma:=\sigma_{v}\otimes_{{\mathcal{O}}}\sigma^{v}$ is an element of ${\mathcal{C}}(K)_{Z}$, then we have a natural isomorphism
$M_{K}(\sigma)\cong\hom_{K_{v}}(M_{K^{v}}(\sigma^{v}),\sigma_{v}^{*})^{*}.$ (22)
4.2. Typicity
$H^{d-1}(\operatorname{Sh}_{K,\bar{\eta}},\overline{\mathbb{Q}}_{l})_{\mathfrak{m}}\cong\bigoplus_{\pi\in{\mathcal{A}}_{K}(\mathfrak{m})}(\pi^{\infty})^{K}\otimes\sigma(\pi^{\infty}),$
where
  • ${\mathcal{A}}_{K}(\mathfrak{m})$ is the set of equivalence classes of automorphic representations of $G({\mathbb{A}})$ with non trivial $K$-invariants and such that its modulo $l$ Satake’s parameters, outside the set $S$ of places dividing the level $K$, are prescribed by $\mathfrak{m}$,
  • and $\sigma(\pi^{\infty})$ is a representation of $\operatorname{Gal}_{F,S}$.
As $\overline{\rho}_{\mathfrak{m}}$ is supposed to be absolutely irreducible, then as explained in chapter VI of [HT01], if $\sigma(\pi^{\infty})$ is non zero, then $\pi$ is a weak transfer of a cohomological automorphic representation $(\Pi,\psi)$ of $\operatorname{GL}_{d}({\mathbb{A}}_{F})\times{\mathbb{A}}_{F}^{\times}$ with $\Pi^{\vee}\cong\Pi^{c}$ where $c$ is the complex conjugation. Attached to such a $\Pi$ is a global Galois representation $\rho_{\Pi,l}:\operatorname{Gal}_{F,S}\longrightarrow\operatorname{GL}_{d}(\overline{\mathbb{Q}}_{l})$ which is irreducible.
Theorem 4.2.1.
(cf. [NF19] theorem 2.20)
If $\rho_{\Pi,l}$ is strongly irreducible, meaning it remains irreducible when it is restricted to any finite index subgroup, then $\sigma(\pi^{\infty})$ is a semi-simple representation of $\operatorname{Gal}_{F,S}$.
Remark. The Tate conjecture predicts that $\sigma(\pi^{\infty})$ is always semi-simple.
Definition 4.2.2.
Proposition 4.2.3.
We suppose that for all $\pi\in{\mathcal{A}}_{K}(\mathfrak{m})$, the Galois representation $\sigma(\pi^{\infty})$ is semi-simple. Then $\mathfrak{m}$ is KHT-typic for $K$.
Proof.
By proposition 5.4 of [Sch18] it suffices to deal with ¯$\overline{\mathbb{Q}}_{l}$-coefficients. From [HT01] proposition VII.1.8 and the semi-simplicity hypothesis, then $\sigma(\pi^{\infty})\cong\widetilde{R}(\pi)^{\bigoplus n(\pi)}$ where $\widetilde{R}(\pi)$ is of dimension $d$. We then write
and $(\pi^{\infty})^{K}\otimes_{\overline{\mathbb{Q}}_{l}}\sigma(\pi^{\infty})\cong((\pi^{\infty})^{K})^{\bigoplus n(\pi)}\otimes_{{\mathbb{T}}(K)_{\mathfrak{m},\overline{\mathbb{Q}}_{l}}}({\mathbb{T}}(K)_{\mathfrak{m},\overline{\mathbb{Q}}_{l}})^{d}$ and finally
In particular with the notations of §4.1, we have an isomorphism
$M_{K^{v}}(\sigma^{v})\otimes_{{\mathbb{T}}(\sigma_{v})_{\mathfrak{m}}}\rho(\sigma^{v})_{\mathfrak{m}}\longrightarrow S_{K^{v}}(\sigma^{v})_{\mathfrak{m}}.$
4.3. Lattices
For each place $w\in S\setminus\{v\}$ choose an inertial type $\tau_{w}$ and a lattice $\sigma^{0}(\tau_{w})$ in $\sigma_{L}(\tau_{w})$ and let $\sigma^{v}:=\otimes_{w\in S\setminus\{v\}}\sigma^{0}(\tau_{w})$ the corresponding representation of $K^{v}=\prod_{w\in S\setminus\{v\}}K_{w}$.
Recall that ${\mathcal{O}}_{L}$ is the ring of integers of a large enough finite extension $L$ of ${\mathbb{Q}}_{l}$. Let $\lambda:{\mathbb{T}}_{\mathfrak{m}}\longrightarrow{\mathcal{O}}_{L}$ be a system [^16] Note that $\lambda\mod\varpi_{L}$ is given by $\mathfrak{m}$, i.e. the modulo $l$ reduction of $\lambda$ is fixed. of Hecke eigenvalues corresponding to some minimal prime ideal of ${\mathbb{T}}_{\mathfrak{m}}:={\mathbb{T}}(\sigma^{v}\sigma^{0}(\tau_{v}))_{\mathfrak{m}}$ for some lattice $\sigma^{0}(\tau_{v})$ of $\sigma_{L}(\tau_{v})$ define just after: note that ${\mathbb{T}}_{\mathfrak{m}}$ does not depend on this choice. Let $\Pi_{v}:=(M_{K^{v}}(\sigma^{v})^{*}[1/l])[\lambda]$ which by strong multiplicity one is an irreducible representation of $\operatorname{GL}_{d}(F_{v})$. We suppose that $\sigma_{L}(\tau_{v})$ appears with multiplicity one in $(\Pi_{v})_{|K_{v}}$. The eigenspace $M_{K^{v}}(\sigma^{v})^{*}[\lambda]$ is a lattice in $\Pi_{v}$ and
$\sigma_{\lambda}(\tau_{v}):=\sigma_{L}(\tau_{v})\cap M_{K^{v}}(\sigma^{v})^{*}[\lambda]$
is $K_{v}$-stable lattice in $\sigma_{L}(\tau_{v})$.
For any lattice $\sigma^{1}(\tau_{v})$ of $\sigma_{L}(\tau_{v})$, it follows from (22) that there is a natural isomorphism
$\hom_{K_{v}}(\sigma^{1}(\tau_{v}),\sigma_{\lambda}(\tau_{v}))^{*}\cong M_{K^{v}}(\sigma^{v}\sigma^{1}(\tau_{v}))/\lambda.$ (23)
Consider some injection
$\iota:\sigma_{1}(\tau_{v})\hookrightarrow\sigma_{2}(\tau_{v})$
between two lattices of $\sigma_{L}(\tau_{v})$. Then $\iota$ induces a injective map
$M_{K^{v}}(\sigma^{v}\sigma_{1}(\tau_{v}))\hookrightarrow M_{K^{v}}(\sigma^{v}\sigma_{2}(\tau_{v})).$
Lemma 4.3.1.
Suppose there exists $\lambda_{0}$ so that the natural map induces by $\iota$
$M_{K^{v}}(\sigma^{v}\sigma_{1}(\tau_{v}))/\lambda_{0}\hookrightarrow M_{K^{v}}(\sigma^{v}\sigma_{2}(\tau_{v}))/\lambda_{0}$ (24)
is an isomorphism, then this is true for every $\lambda\equiv\lambda_{0}\mod\varpi_{L}$.
Proof.
For any $\lambda$ the natural inclusion
${\mathcal{O}}_{L}\cong M_{K^{v}}(\sigma^{v}\sigma_{1}(\tau_{v}))/\lambda\hookrightarrow M_{K^{v}}(\sigma^{v}\sigma_{2}(\tau_{v}))/\lambda\cong{\mathcal{O}}_{L}$
is either an isomorphism or multiplication by some power of $\varpi_{L}$ and so zero modulo $\varpi_{L}$. By hypothesis modulo $\varpi_{L}$, the natural map $\iota$ induces an isomorphism
$M_{K^{v}}(\sigma^{v}\sigma_{1}(\tau_{v}))/(\lambda_{0},\varpi_{L})\cong M_{K^{v}}(\sigma^{v}\sigma_{2}(\tau_{v}))/(\lambda_{0},\varpi_{L}).$
As the following diagram is commutative
[…]
with common kernel $\mathfrak{m}$, we then deduce that $\iota$ induces an isomorphism
$M_{K^{v}}(\sigma^{v}\sigma_{1}(\tau_{v}))/(\lambda,\varpi_{L})\cong M_{K^{v}}(\sigma^{v}\sigma_{2}(\tau_{v}))/(\lambda,\varpi_{L}),$
so that the natural map induces by $\iota$
$M_{K^{v}}(\sigma^{v}\sigma_{1}(\tau_{v}))/\lambda\hookrightarrow M_{K^{v}}(\sigma^{v}\sigma_{2}(\tau_{v}))/\lambda$
is also an isomorphism. ∎
Proposition 4.3.2.
Consider two automorphic irreducible representations $\Pi$ and $\Pi^{\prime}$ associated to respectively two systems171717with modulo $l$ reduction fixed by $\mathfrak{m}$ of Hecke eigenvalues $\lambda,\lambda^{\prime}:{\mathbb{T}}_{\mathfrak{m}}\longrightarrow{\mathcal{O}}_{L}$. We denote by
$\Pi_{v}:=(M_{K^{v}}(\sigma^{v})^{*}[1/l])[\lambda]\quad\hbox{and}\quad\Pi^{\prime}_{v}:=(M_{K^{v}}(\sigma^{v})^{*}[1/l])[\lambda^{\prime}].$
Consider a $K$-type $\tau_{v}$ such that $\sigma_{L}(\tau_{v})$ appears with multiplicity one in $(\Pi_{v})_{|K_{v}}$ (resp. in $(\Pi^{\prime}_{v})_{|K_{v}}$) and let $\sigma_{\lambda}(\tau_{v})$ (resp. $\sigma_{\lambda^{\prime}}(\tau_{v})$) be the stable lattice in $\sigma_{L}(\tau_{v})$ defined by
$\sigma_{\lambda}(\tau_{v}):=\sigma_{L}(\tau_{v})\cap M_{K^{v}}(\sigma^{v})^{*}[\lambda],\quad\hbox{resp. }\sigma_{\lambda^{\prime}}(\tau_{v}):=\sigma_{L}(\tau_{v})\cap M_{K^{v}}(\sigma^{v})^{*}[\lambda^{\prime}].$
Then these two lattices are homothetic.
Proof.
Up to homothety any stable lattice $\sigma_{+}$ is such that
$\sigma_{0}:=\sigma_{\lambda}(\tau_{v})\subsetneq\sigma_{+}\subsetneq\varpi_{L}^{-\delta}\sigma_{\lambda}(\tau_{v}),$
where $\delta$ is minimal and $\sigma_{0}\not\subseteq\varpi_{L}\sigma_{+}$. The natural morphism
${\mathcal{O}}_{L}\cong\hom_{\operatorname{GL}_{d}({\mathcal{O}}_{v})}(\sigma_{+},\sigma_{\lambda}(\tau_{v}))\longrightarrow\hom_{\operatorname{GL}_{d}({\mathcal{O}}_{v})}(\sigma_{0},\sigma_{\lambda}(\tau_{v}))\cong{\mathcal{O}}_{L}$
is given by multiplication by $\varpi_{L}^{\delta}$, while
$\hom_{\operatorname{GL}_{d}({\mathcal{O}}_{v})}(\sigma_{0},\sigma_{\lambda}(\tau_{v}))\longrightarrow\hom_{\operatorname{GL}_{d}({\mathcal{O}}_{v})}(\sigma_{-},\sigma_{\lambda}(\tau_{v}))$
is a isomorphism. From the isomorphism (23), $\iota$ induces
$M_{K^{v}}(\sigma^{v}\sigma_{-})/\lambda=M_{K^{v}}(\sigma^{v}\sigma_{0})/\lambda=\varpi_{L}^{\delta}M_{K^{v}}(\sigma^{v}\sigma_{+})/\lambda.$ (25)
If $\sigma_{\lambda^{\prime}}(\tau_{v})$ were equal to $\sigma_{+}$ then the morphism induced by $\iota$
$\hom_{\operatorname{GL}_{d}({\mathcal{O}}_{v})}(\sigma_{+},\sigma_{\lambda^{\prime}}(\tau_{v}))\hookrightarrow\hom_{\operatorname{GL}_{d}({\mathcal{O}}_{v})}(\sigma_{0},\sigma_{\lambda^{\prime}}(\tau_{v}))$
would be an isomorphism and by (23), the natural injection would give the equality
$M_{K^{v}}(\sigma^{v}\sigma_{0})/\lambda^{\prime}=M_{K^{v}}(\sigma^{v}\sigma_{+})/\lambda^{\prime},$
which, by the previous lemma, is not compatible with (25). ∎
5. Proof of Ihara’s lemma
@@ 5.2. Level raising
Fact from [Boy10] §3: for an irreducible cuspidal representation $\pi_{v}$ such that its modulo $l$ reduction is isomorphic to $\varrho$, $H^{0}(\operatorname{Sh}_{K^{v}(\infty),\bar{s}_{v}},P(t,\pi_{v}))_{\mathfrak{m}}\otimes_{\overline{\mathbb{Z}}_{l}}\overline{\mathbb{Q}}_{l}$ is the sum of the contributions of $\Pi_{\widetilde{\mathfrak{m}}}$ with $\widetilde{\mathfrak{m}}\subseteq\mathfrak{m}$ such that $\Pi_{\widetilde{\mathfrak{m}}}$ is of the following shape: $\mathrm{st}_{t}(\pi^{\prime}_{v})\times\psi$ where $\pi^{\prime}_{v}$ is an unramified twist of $\pi_{v}$ and $\psi$ is any representation of $\operatorname{GL}_{d-tg}(F_{v})$ whose cuspidal support is not linked to those of $\mathrm{st}_{t}(\pi^{\prime}_{v})$.